# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v1

@@ Abstract
Recent generative models based on score matching and flow matching have significantly advanced generation tasks, but their potential in discriminative tasks remains underexplored. Previous approaches, such as generative classifiers, have not fully leveraged the capabilities of these models for discriminative tasks due to their intricate designs. We propose Pretrained Reversible Generation (PRG), which extracts unsupervised representations by reversing the generative process of a pretrained continuous flow model. PRG effectively reuses unsupervised generative models, leveraging their high capacity to serve as robust and generalizable feature extractors for downstream tasks. Our method consistently outperforms prior approaches across multiple benchmarks, achieving state-of-the-art performance among generative model-based methods, including 78% top-1 accuracy on ImageNet. Extensive ablation studies further validate the effectiveness of our approach.
@@ 1 Introduction
Figure 1: Overview of the PRG as Unsupervised Visual Representation pipeline. Swiss-roll data is generated using $(x,y)=\left(t\cos(t),t\sin(t)\right)$ with $t\in[0,3\pi]$, where the color gradient from blue to red corresponds to $t$ increasing from 00 to $3\pi$. The yellow trajectory represents the generative process, with color intensity indicating direction. Green arrows depict the reverse generation process; Using the pretrained model, each step along the green path can be fine-tuned for downstream tasks.
@@
Generative models formulated as continuous-time stochastic differential equations, including diffusion and flow models [lipman2023flow, liu2022flow, pmlr-v202-pooladian23a, Albergo2023], have shown remarkable proficiency in multi-modal content generation tasks [rombach2022high, Ho2022video, mittal2021symbolic], as well as in scientific modeling [abramson2024accurate]. By effectively learning high-dimensional distributions, these models can generate new, high-quality samples that resemble the original data.
@@
While generative tasks necessitate models to reconstruct the entire data distribution, another category, known as discriminative tasks (e.g., image classification), involve learning representations that capture the essential underlying data structure and output discrete or continuous predictions based on learned representations. Previous works have attempted to enhance representation learning through various methods, with particular attention on unsupervised learning recently. These methods endeavor to construct generic proxy tasks that effectively acquire and leverage data for pretraining, thereby providing robust representations for different downstream tasks. [bengio2013representation, radford2015unsupervised, chen2016infogan, van2017neural, donahue2019large, chen2020generative].
@@
Due to the extraordinary performance of diffusion models in generative tasks, recent studies have explored their potential for discriminative tasks [clark2024text, li2023your]. These studies often rely on extracting or combining representations from the internal layers of pretrained diffusion models, involving intricate and non-generalizable designs, yet usually lacking a clear explanation. Moreover, the performance of these methods exhibits significant room for improvement compared to current state-of-the-art techniques.
Inspired by the reversible property of continuous-time stochastic flow models, in this paper, we introduce a general, model-agnostic, and effective method to apply generative models to discriminative tasks by reversing the generative process into a meaningful inference process. This method does not require accessing any internal features of a pretrained generative model, thus avoiding the need for fixed architectural modules to capture data representations. Another advantage of this method is that it enables unsupervised representation learning, making it applicable to unsupervised learning tasks without additional cost.
We demonstrate that this method is effective across diverse generative models. Through theoretical analysis and empirical experiments, we explain why reversing a pretrained generative process can serve as an effective feature extractor for downstream tasks. Our experiments show that this method achieves competitive performance in both supervised and unsupervised image classification tasks.
@@
  • Extensive experiments demonstrate the robustness and generalizability of our method on several image classification datasets.
2 Preliminaries
We introduce the generative models and mathematical notations used in this work, focusing on training and inference within diffusion and flow models.
2.1 Continuous-time Stochastic Flow Models
We consider two continuous-time stochastic flow generative models: diffusion models and flow models. Let $t$ represent time, with the interval $[1,0]$ for the generation process and $[0,1]$ for the reverse of generation process. We denote the data and its transformations as $x_{t}$, where $x_{0}$ is the original data from the dataset and $x_{1}$ is the terminal state, achieved by reversing the generative process.
2.1.1 Diffusion Model
A diffusion model reverses a forward diffusion process governed by the following SDE:
$\mathrm{d}x=f(t,x_{t})\mathrm{d}t+g(t)\mathrm{d}w_{t},$ (1)
where $f(t,x_{t})$ is the drift term, $g(t)$ is the diffusion coefficient, and $w_{t}$ is the Wiener process.
For simplicity, we use an SDE with a linear drift:
$\mathrm{d}x=f(t)x_{t}\mathrm{d}t+g(t)\mathrm{d}w_{t}.$ (2)
The transition distribution of $x\in\mathbb{R}^{d}$ from time 00 to $t$ is given by:
$p(x_{t}|x_{0})\sim\mathcal{N}(x_{t}|\alpha_{t}x_{0},\sigma_{t}^{2}I),$ (3)
where $\alpha_{t}$ is a scaling factor and $\sigma_{t}^{2}$ is the variance.
The diffusion process is predefined, such as the Variance-Preserving SDE (VP-SDE) [song2020score] or the Generalized VP-SDE (GVP) [Albergo2023]. Full definitions are provided in the Appendix A. The marginal distribution $p(x_{t})$ of the diffusion process is fixed if the data distribution $p(x_{0})$ is stationary. Using the Fokker-Planck equation, an equivalent ODE can be derived with the same marginal distribution, known as the Probability Flow ODE [song2020score]:
In practice, the velocity model $v(x_{t})$ or score function $\nabla_{x_{t}}\log p(x_{t})$ in Eq. 4 is parameterized as $v_{\theta}$ or $s_{\theta}$ using a neural network, such as in Diffusion Transformers [peebles2023scalable, ma2024sit].
2.1.2 Flow Model
A flow model defines a deterministic transformation from $x_{0}$ to $x_{1}$ via a parameterized velocity function $v(x_{t})$, governed by the ODE: $\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=v(x_{t})$. The marginal probability distribution $p(x_{t})$ evolves through the vector field $v(x_{t})$, following the continuity equation:
$\frac{\partial p(x_{t})}{\partial t}+\nabla_{x}\cdot(p(x_{t})v(x_{t}))=0,$ (5)
which transports $p(x_{0})$ to $p(x_{1})$ along the flow path.
For the Independent Conditional Flow Matching (ICFM) models [tong2024improving] considered in this paper, the flow path is predefined as $p(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t}|tx_{1}+(1-t)x_{0},\sigma^{2}I)$, with the velocity function $v(x_{t}|x_{0},x_{1})=x_{1}-x_{0}$. We also studied Optimal Transport Conditional Flow Matching (OTCFM) models [tong2024improving]. The marginal distribution $p(x_{t})$ remains stationary if both $p(x_{0})$ and $p(x_{1})$ are stationary.
2.2 Generative Model Training and Inference
Both the diffusion and flow models can be parameterized by $v_{\theta}(x_{t})$ and pretrained using the Conditional Flow Matching loss, denoted $\mathcal{L}_{\text{FM}}$ [lipman2022flow, tong2024improving]. In the vanilla case, the weighting factor $\lambda_{\mathrm{FM}}$ equals $1$. For diffusion models, score matching is also applicable, with $\lambda_{\mathrm{SM}}(t)=\sigma_{t}^{2}$ in the vanilla case.
$\mathcal{L}_{\text{FM}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[\lambda_{\mathrm{FM}}(t)\|v_{\theta}(x_{t})-v(x_{t})\|^{2}\right]\mathrm{d}t,$ (6)
$\mathcal{L}_{\text{SM}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[\lambda_{\mathrm{SM}}(t)\|s_{\theta}(x_{t})-\nabla_{x_{t}}\log p(x_{t})\|^{2}\right]\mathrm{d}t.$
Once trained, starting with $x_{1}$ sampled from a standard Gaussian, we can recover $x_{0}$ from $x_{1}$ or infer the latent variable $x_{1}$ from $x_{0}$ by reversing the generative process.
@@ 3 Related Work
Generative classifiers is first proposed in [li2023your, chen2023robust, clark2024text, chendiffusion, zimmermann2021score, pmlr-v162-nie22a] model the conditional distribution $p(x|y)$ and use Bayes’ rule to compute the probability $p(y|x)=\frac{p(x|y)p(y)}{p(x)}$. These classifiers are typically trained by maximizing the conditional log-likelihood: $\max_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}\log p(x|y)$.
When employing conditional diffusion models, $p(x|y)$ can be computed with the instantaneous change-of-variables formula [Chen2018, Grathwohl2018] or approximated via ELBO [sohl2015deep, ho2020denoising]. duvenaud2020your advocate an energy-based model to enhance calibration, robustness, and out-of-distribution detection. yang2022your propose HybViT, which combines diffusion models and discriminative classifiers to predict $p(y|x)$. zimmermann2021score introduce score-based generative classifiers but find them vulnerable to adversarial attacks. clark2024text and li2023your explore zero-shot generative classifiers using pretrained models like Stable Diffusion [rombach2022high] to denoise images based on text labels.
In contrast, our approach fine-tunes a classifier on top of the generative model, making it a discriminative classifier that directly estimates $p(y|x)$. Discriminative classifiers are often preferred for classification tasks [Vapnik1995, AndrewNIPS2001]. fetaya2019understanding suggest a trade-off between likelihood-based modeling and classification accuracy, implying that improving generation may hinder classification for generative classifier. However, our two-stage training scheme shows that enhancing the generative model in the pretraining stage can improve discriminative performance during the finetuning stage. This suggests a complementary relationship between likelihood-based training and classification accuracy, indicating that discriminative models can benefit from generative pretraining. Thus, we demonstrate a viable approach to leveraging generative models for discriminative tasks.
@@
Generative models have been widely explored for visual representation learning [kingma2013auto, van2017neural, goodfellow2020generative]. Denoising autoencoders (DAEs) [vincent2008extracting, vincent2010stacked, bengio2013generalized, geras2014scheduled] learn robust representations by reconstructing corrupted input data, which can be interpreted through a manifold learning perspective [vincent2011]. Our method extends vincent2010stacked by leveraging reversible flow-based generative models like diffusion models.
Diffusion models [sohl2015deep, ho2020denoising, song2021maximum] have shown promise in generation, yet their potential for feature extraction is underexplored. chen2020generative pretrain a Transformer for autoregressive pixel prediction, demonstrating competitive results in representation learning. Their two-stage approach—pretraining a generative model followed by full model finetuning—bears similarity to ours. Additionally, their downsampling of images to low resolution mimics the effect of adding noise, analogous to denoising. However, their model and autoregressive generation scheme are unsuitable for reverse generation and lack efficiency. he2022masked pretrain a masked autoencoder, effectively performing denoising by reconstructing incomplete data. baranchuk2021label show diffusion models improve semantic segmentation, particularly with limited labeled data. sinha2021dc propose Diffusion-Decoding with contrastive learning to enhance representation quality, while mittal2023diffusion introduce a Diffusion-based Representation Learning (DRL) framework, leveraging denoising score matching.
Recently, xiang2023denoising investigate using denoising diffusion autoencoders for discriminative tasks via generative pre-training. Our method differs by using the reverse generative process to extract features, offering a more general and model-agnostic approach.
@@ 4 Method
To address the issues discussed in Section 3, We propose an approach follows the conventional two-stage training scheme commonly used in generative modeling applications. In the first stage, we pretrain diffusion or flow models in an unsupervised manner. In the second stage, we fine-tune the models for discriminative tasks by conducting inference in the reverse direction of the generative process.
4.1 Pretraining Mechanism
@@
In the traditional autoencoder paradigm, representation learning aims to learn a network that maps the input data $X$ into a useful representation $Z$ for downstream tasks. We parameterize the encoder $p_{\theta}(z|x)$ with parameters $\theta$. A good representation should retain sufficient information about the input $X$. In information-theoretic terms [linsker1988application, vincent2010stacked], this corresponds to maximizing the mutual information $\mathcal{I}(X,Z)$ between the input random variable $X$ and its representation $Z$. Following the reasoning of vincent2010stacked, we show that pretraining diffusion or flow models maximizes the mutual information. Specifically, we aim to find the optimal $\theta^{*}$ that maximizes $\mathcal{I}(X,Z)$:
$\theta^{*}$ $=\arg\max_{\theta}\mathcal{I}(X,Z)$ (7)
$=\arg\max_{\theta}[\mathcal{H}(X)-\mathcal{H}(X|Z)]$
$=\arg\max_{\theta}[-\mathcal{H}(X|Z)]$
$=\arg\max_{\theta}\mathbb{E}_{p(z,x)}[\log p(x|z)].$
@@
In continuous-time stochastic flow generative models, the inference and generation process are reversible via a same model, which means the decoder parameter $\theta^{\prime}$ and encoder parameter $\theta$ are shared parameters. However, we still use these two notation in the later part of this paper for distinguishing the inference and generation process through ODE solver with $\theta$ and $\theta^{\prime}$ respectively. For preventing any confusion, we denote $x_{0}$ being data variable $x$ amd $x_{1}$ being the inferred latent variable $z$. In practice, training generative models using score matching or flow matching objectives does not directly maximize the likelihood of reconstructing $x_{0}$ from $x_{1}$. However, as shown in previous works [song2021maximum, lu2022maximum, zheng2023improved], these methods provide a lower bound on the likelihood. Specifically, they establish that:
@@
From Eq. 11, we observe that using score matching objectives during training can help increase the data likelihood. Flow matching differs from score matching by employing a slightly different weighting factor in the loss function, so it can also be used to increase the likelihood, as shown in Eq. 6. Therefore, pretraining diffusion or flow models via score matching or flow matching can be viewed as maximizing the mutual information between the data and its representation encoded through the reverse generative process. This relationship between the mutual information and the training loss is guaranteed by the Eq. 11. This process transforms the data distribution into a Gaussian distribution while preserving as much information as possible.
4.2 Finetuning by Reversing Generation
Due to the invertibility of the generative process (Eq. 4), the latent variable $x_{1}$ corresponding to a data point $x_{0}$ can be obtained by reversing the generative process. We denote $x_{t}=F_{\theta}(x_{0})$ as the extracted features for downstream discriminative tasks, such as classification or regression.
This paper focuses on supervised learning for downstream tasks. Suppose we have a labeled dataset $\mathcal{D}_{\text{finetune}}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is a data point and $y_{i}$ is its label, potentially differing from the pretraining dataset $\mathcal{D}_{\text{pretrain}}$. We introduce a classifier $p_{\phi}(y\mid z)$, parameterized by $\phi$, trained by minimizing the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}\log p_{\phi}(y_{i}\mid F_{\theta}(x_{i})).$ (13)
Finetuning Strategy
Generative features may not be optimal for discriminative tasks, as reconstruction requires more details than classification. Hence, fine-tuning is necessary to adapt the model for downstream tasks.
A key question is whether the flow model should also be fine-tuned alongside the classifier. Due to the flow model’s large capacity, reusing the pretrained model seems preferable. Initially, we explored freezing the flow model while finetuning only the classifier. However, this approach underperformed, indicating that fine-tuning both the flow model and the classifier is necessary. Otherwise, a large classifier with extra parameters would be required, reducing efficiency. To fine-tune the entire model, we leverage neural ODEs [Chen2018] for efficient gradient computation, optimizing both $\theta$ and $\phi$ via backpropagation.
We observe that the generative features are often redundant for discriminative tasks and are pruned during fine-tuning, as we will empirically demonstrate in the experiments section. Additionally, we assess the robustness and generalization of these features against out-of-distribution samples through further experiments.
Classifier Design
We found that the features $x_{t}$ are highly linearly separable, allowing a simple neural network to perform well without compromising accuracy. In practice, we also apply label smoothing [muller2019does] to enhance generalization. In this paper, we use a two-layer MLP with tanh activations as an example. Notably, increasing the classifier’s size did not significantly improve performance, suggesting that the features extracted by the reversed generative process are already highly informative.
Figure 2: CIFAR-10 Accuracy (y-axis) after fine-tuning checkpoints from different pretraining stages (x-axis).
Figure 3: Evaluation of the indicator (y-axis) $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}|x_{1})]$ from different pretraining stages (x-axis).

### v2

@@ Abstract
Recent generative models based on score matching and flow matching have significantly advanced generation tasks, but their potential in discriminative tasks remains underexplored. Previous approaches, such as generative classifiers, have not fully leveraged the capabilities of these models for discriminative tasks due to their intricate designs. We propose Pretrained Reversible Generation (PRG), which extracts unsupervised representations by reversing the generative process of a pretrained continuous generation model. PRG effectively reuses unsupervised generative models, leveraging their high capacity to serve as robust and generalizable feature extractors for downstream tasks. This framework enables the flexible selection of feature hierarchies tailored to specific downstream tasks. Our method consistently outperforms prior approaches across multiple benchmarks, achieving state-of-the-art performance among generative model based methods, including 78% top-1 accuracy on ImageNet at a resolution of 64×64. Extensive ablation studies, including out-of-distribution evaluations, further validate the effectiveness of our approach.
@@ 1 Introduction
Figure 1: Overview of the PRG as Unsupervised Visual Representation pipeline. Swiss-roll data is generated using $(x,y)=\left(t\cos(t),t\sin(t)\right)$ with $t\in[0,3\pi]$, where the color gradient from blue to red corresponds to $t$ increasing from 00 to $3\pi$. The yellow trajectory represents the generative (forward) process, with color intensity indicating direction. Green arrows illustrate the inference (reverse) generation process. Using the pretrained model, each step along the green path can be fine-tuned for downstream tasks.
@@
Generative models formulated as continuous-time stochastic differential equations, including diffusion and flow models [36, 38, 50, 2], have shown remarkable proficiency in multi-modal content generation tasks [52, 27, 44], as well as in scientific modeling [1]. By effectively learning high-dimensional distributions, these models can generate new, high-quality samples that resemble the original data.
@@
While generative tasks require models to reconstruct the entire data distribution, discriminative tasks (e.g., image classification) rely on learning representations where some parts capture the essential underlying data structure to produce discrete or continuous predictions. Previous works [17, 23] have explored various approaches to enhance representation learning, with growing focus on unsupervised visual representation learning [12, 10]. These methods construct generic proxy tasks to efficiently extract and utilize data for pre-training, yielding robust representations for diverse downstream tasks [5, 51, 13, 59, 16, 10, 65].
@@
Due to the extraordinary performance of diffusion models in generative tasks, recent studies have explored their potential for discriminative tasks [14, 34]. These studies often rely on extracting or combining representations from the internal layers of pretrained diffusion models, involving intricate and poorly generalizable designs, yet usually lacking a clear explanation. For example, [65] selects a specific layer in a UNet-based diffusion model as the feature representation. Moreover, the performance of these methods exhibits significant room for improvement compared to current state-of-the-art techniques.
Inspired by the reversible property [11, 56] of continuous-time stochastic flow models, we introduce Pretrained Reversible Generation as Unsupervised Visual Representation Learning (PRG). Our approach applies generative models to discriminative tasks by reversing the generative process into a meaningful inference process. Specifically, we pretrain a reversible generative model via flow matching to maximize the lower bound of mutual information between the original image and its optimal representation, providing a strong initialization. We then reverse the generative process and leverage it for downstream training to improve task performance. This model-agnostic method does not require access to any internal features of a pretrained generative model, eliminating the need for fixed network modules to capture data representations. As a pretrain-finetune paradigm, it fully leverages the powerful capabilities of various pretrained generative models (e.g., diffusion models, flow models). During the reversed process, it utilizes features from different hierarchical levels, enabling efficient adaptation to downstream tasks. Through theoretical analysis Sec. 3.2 and verification experiments Sec. 4.3, we explain why reversing a pretrained generative process serves as an effective feature extractor for downstream tasks. Our main experiments further demonstrate that this method achieves competitive performance in image classification and OOD detection while effectively adapting community-pretrained text-to-image models for downstream training.
@@
  • Extensive experiments demonstrate the robustness and generalizability of PRG in many image classification datasets.
@@ 2 Related Work
Generative classifiers is first proposed in [34, 9, 14, 8, 73, 49] model the conditional distribution $p(x|y)$ and use Bayes’ rule to compute the probability $p(y|x)=\frac{p(x|y)p(y)}{p(x)}$. These classifiers are typically trained by maximizing the conditional logarithmic likelihood: $\max_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}\log p(x|y)$.
When using conditional diffusion models, $p(x|y)$ can be computed with the instantaneous change-of-variable formula [11, 22] or approximated by ELBO [55, 26]. Duvenaud et al. [18] advocate an energy-based model to improve calibration, robustness, and out-of-distribution detection. Yang et al. [67] propose HybViT, which combines diffusion models and discriminative classifiers to predict $p(y|x)$. Zimmermann et al. [73] introduce score-based generative classifiers but find them vulnerable to adversarial attacks. Clark and Jaini [14] and Li et al. [34] explore zero-shot generative classifiers using pretrained models like Stable Diffusion [52] to denoise images based on text labels.
In contrast, our approach fine-tunes a classifier on top of the generative model, making it a discriminative classifier that directly estimates $p(y|x)$. Discriminative classifiers are often preferred for classification tasks [60, 47]. Fetaya et al. [19] suggest a trade-off between likelihood-based modeling and classification accuracy, implying that improving generation may hinder classification for generative classifier. However, our two-stage training scheme shows that enhancing the generative model in the pretraining stage can improve discriminative performance during the fine-tuning stage. This suggests a complementary relationship between likelihood-based training and classification accuracy, indicating that discriminative models can benefit from generative pretraining. Thus, we demonstrate a viable approach to leveraging generative models for discriminative tasks.
@@
Generative models have been widely explored for visual representation learning [29, 59, 21]. Denoising autoencoders (DAEs) [62, 63, 6, 20] learn robust representations by reconstructing corrupted input data, which can be interpreted through a manifold learning perspective [61]. Our method extends Vincent et al. [63] by leveraging reversible flow-based generative models like diffusion models.
Diffusion models [55, 26, 57] have shown promise in generation, yet their potential for feature extraction is underexplored. Chen et al. [10] pretrain a Transformer for autoregressive pixel prediction, demonstrating competitive results in representation learning. Their two-stage approach (pre-training a generative model followed by full fine-tuning) closely resembles ours. Additionally, their downsampling of images to low resolution mimics the effect of adding noise, analogous to denoising. However, their model and autoregressive generation scheme are unsuitable for reverse generation and lack efficiency. He et al. [24] pretrain a masked autoencoder, effectively performing denoising by reconstructing incomplete data. Baranchuk et al. [4] show diffusion models improve semantic segmentation, particularly with limited labeled data. Sinha et al. [54] propose Diffusion-Decoding with contrastive learning to enhance representation quality, while Mittal et al. [45] introduce a Diffusion-based Representation Learning (DRL) framework, leveraging denoising score matching.
Compared to previous methods, our approach demonstrates that a pretraining-finetuning pipeline enhances discriminative performance, indicating complementarity rather than a tradeoff. Unlike strategies that refine diffusion training or layer architectures, we leverage diffusion models to build an ODE-based classifier, achieving outstanding downstream results across multiple datasets.
@@ 3 Method
To address the issues discussed in Section 2, We propose an approach that follows the conventional two-stage training scheme commonly used in generative modeling applications. In the first stage, we pretrain diffusion or flow models in an unsupervised manner. In the second stage, we fine-tune the models for discriminative tasks by conducting inference in the reverse direction of the generative process.
3.1 Pretrained Reversible Generation Framework
We consider two continuous-time stochastic flow generative models: diffusion and flow models. Let $t$ denote time, where $t\in[1,0]$ corresponds to the generation, and $t\in[0,1]$ to its reverse. We denote the data and their transformations as $x_{t}$, with $x_{0}$ as the original and $x_{1}$ as the terminal state.
A diffusion model reverses a forward diffusion process described by $p(x_{t}\mid x_{0})\sim\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}I)$, where $\alpha_{t}$ is a scaling factor and $\sigma_{t}^{2}$ denotes the variance. We adopt the Generalized VP-SDE, which uses a triangular function as coefficients to define the diffusion path, referred to as PRG-GVP with $\alpha_{t}=\cos\left(\frac{1}{2}\pi t\right)$ and $\sigma_{t}=\sin\left(\frac{1}{2}\pi t\right)$. Due to the Fokker-Planck equation, an equivalent ODE can be derived that shares the same marginal distribution, known as the Probability Flow ODE [56]:
A flow model deterministically transforms $x_{0}$ to $x_{1}$ via a parameterized velocity function $v(x_{t})$. Inspired by [58], we adopt the flow path formulation, referred to as PRG-ICFM, with $p(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t}|tx_{1}+(1-t)x_{0})$, and velocity function: $v(x_{t}|x_{0},x_{1})=x_{1}-x_{0}$.
We also employ the Optimal Transport Conditional Flow Matching (OTCFM) model [58]. Specifically, we jointly sample $(x_{0},x_{1})$ from the optimal transport plan $\pi$ to construct the 2-Wasserstein target distribution. The corresponding velocity field $v(x_{t}\mid x_{0},x_{1})$ is then used to approximate dynamic optimal transport, referred to as PRG-OTCFM. In practice, we parameterize the velocity model $v(x_{t})$ as $v_{\theta}$ using a network to model the above variants of PRG.
$\mathcal{L}_{\text{FM}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[\lambda_{\mathrm{FM}}(t)\|v_{\theta}(x_{t})-v(x_{t})\|^{2}\right]\mathrm{d}t$ (2)
After completing pre-training via Eq. 2, and given the invertibility of Eq. 1, we obtain $x_{1}$ for a data point $x_{0}$ through the same reversal. We denote $x_{t}=F_{\theta}(x_{0})$ as the extracted features for downstream discriminative tasks, such as classification or regression.
This paper focuses on supervised learning for downstream tasks while simultaneously improving the lower bound of mutual information. Suppose we have a labeled dataset $\mathcal{D}_{\text{finetune}}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is a data point and $y_{i}$ is its label, potentially differing from the pre-training dataset $\mathcal{D}_{\text{pretrain}}$. Thus, during the fine-tuning stage, we introduce a classifier $p_{\phi}(y\mid z)$, parameterized by $\phi$, minimizing the cross-entropy loss and flow-matching loss:
$\mathcal{L}_{\text{total}}=-\sum_{i=1}^{N}\log p_{\phi}(y_{i}\mid F_{\theta}(x_{i}))+\beta\mathcal{L}_{\text{FM}}(x).$ (3)
Classifier Design
We found that the features $x_{t}$ are well-structured, allowing a simple two-layer MLP with tanh activations to perform well. Notably, enlarging the classifier yielded minimal gains, suggesting that the reversed generative process already captures highly informative features.
Diffusion Step Design
To reduce computational costs, the representations used for inference do not necessarily need to traverse the entire process. Instead, our framework allows fine-tuning and inference to start from any selected point along the trajectory. Moreover, more complex downstream tasks typically require greater training/inference steps. Further details are provided in Secs. 4.3.2 and C.5.
3.2 Pretrained Reversible Generation Mechanism
pre-training
@@ 3.2 Pretrained Reversible Generation Mechanism
In the traditional autoencoder paradigm, representation learning aims to learn a network that maps the input data $X$ into a useful representation $Z$ for downstream tasks. We parameterize the encoder $p_{\theta}(z|x)$ with parameters $\theta$. A good representation should retain sufficient information about the input $X$. In information-theoretic terms [35, 63], this corresponds to maximizing the mutual information $\mathcal{I}(X,Z)$ between the input random variable $X$ and its representation $Z$. Following the reasoning of Vincent et al. [63], we show that pre-training diffusion or flow models maximizes the mutual information. Specifically, we aim to find the optimal $\theta^{*}$ that maximizes $\mathcal{I}(X,Z)$:
$\theta^{*}$ $=\arg\max_{\theta}\mathcal{I}(X,Z)=\arg\max_{\theta}[\mathcal{H}(X)-\mathcal{H}(X|Z)]$ (4)
$=\arg\max_{\theta}[-\mathcal{H}(X|Z)]=\arg\max_{\theta}\mathbb{E}_{p(z,x)}[\log p(x|z)].$
@@
In continuous-time stochastic flow generative models, the inference and generation process are reversible via a same model, which means the decoder parameter $\theta^{\prime}$ and encoder parameter $\theta$ are shared parameters. However, we still use these two notation in the later part of this paper for distinguishing the inference and generation process through ODE solver with $\theta$ and $\theta^{\prime}$ respectively. For preventing any confusion, we denote $x_{0}$ being data variable $x$ amd $x_{1}$ being the inferred latent variable $z$. In practice, training generative models using score matching or flow matching objectives does not directly maximize the likelihood of reconstructing $x_{0}$ from $x_{1}$. However, as shown in previous works [57, 41, 72], these methods provide a lower bound on the likelihood. Specifically, they establish that:
@@
From Eq. 8, we observe that using score matching objectives during training can help increase the data likelihood. Flow matching differs from score matching by employing a slightly different weighting factor in the loss function, so it can also be used to increase the likelihood, as shown in Eq. 2. Therefore, pre-training diffusion or flow models via score matching or flow matching can be viewed as maximizing the mutual information between the data and its representation encoded through the reverse generative process. This relationship between the mutual information and the training loss is guaranteed by the Eq. 8. This process transforms the data distribution into a Gaussian distribution while preserving as much information as possible.
Fine-tuning by Reversing Generation
Generative features may not be optimal for discriminative tasks, as reconstruction requires more detail than comprehension. Hence, fine-tuning is necessary to adapt the model for downstream tasks. A key question is whether the flow model should also be fine-tuned alongside the classifier. Due to the flow model’s large capacity, reusing the pretrained model seems preferable. Initially, we explored freezing the model while fine-tuning only the classifier. However, this approach underperformed, indicating that fine-tuning both the flow model and the classifier is necessary. Otherwise, a large and complex classifier would be required, reducing efficiency. To fine-tune the entire model, we leverage neural ODEs [11] for efficient gradient computation, optimizing both $\theta$ and $\phi$ via backpropagation. We observed that distinct discriminative tasks correspond to different parts of generative features Appendix D. Thus, fine-tuning can selectively prune or enhance relevant features as needed. In addition, we assess the robustness and generalization of these features against out-of-distribution samples through further experiments.
Figure 2: CIFAR-10 Accuracy (y-axis) after fine-tuning checkpoints from different pre-training stages (x-axis).
Figure 3: Evaluation of the indicator (y-axis) $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}|x_{1})]$ from different pre-training stages (x-axis).